40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4155     Entry time: Fri Jan 14 12:29:57 2011
Author: Koji 
Type: Update 
Category: Locking 
Subject: Next steps for the green 

These are the next steps for a better operation of the arm locking. They are suitable for the day time activities

Reconfiguration of the X-End table

- End transmission power monitor (CDS model exists, need to configure the PD)

- IR steering mirror for the transmon

- Restore/align end green beam

- Relocate the end trans CCD

- Connect the video output cable for the X-end CRT monitor

LSC Whitening

- LSC Whitening binary IO connection

 


They are not urgent but also good things to do

MC servo characterization

- Error signal: frequency noise

- Loop characterization

Arm cavity characterization with cavity sweep

- Arm finesse for 1064nm and 532nm

- Arm FSR measurement with 1064 (and optionally with 532nm simultaneously)

- Arm g-factor for 1064nm and 532nm

ELOG V3.1.3-